Can some please help me solve this math problem: Solve the equation 2x^2-5x+2=0 for x. I really appreciate it.
Thank you

Answer by tutorcecilia(2152)

2x^2-5x+2=0
(2x-1)(x-2) [factor]
.
2x-1=0 [set each factor equal to zero and solve for the x-term]
2x=1
x=1/2
.
x-2=0
x=2